# parskip=half

### v2

@@ 2.2.3 Augmented corrections
The main disadvantage of the two previous transformations is that they heavily rely on the specification of good estimator $\widehat{G}$ (for IPCW) or $\widehat{S}$ (for BJ). In order to circumvent this limitation, Rubin and van der Laan, (2007) proposed the following transformations, whose expression is based on theory of semi-parametric estimation developed in Van der Laan and Robins, (2003),
@@
where $\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A)$ is the distribution of $C$ conditional on the covariates $X$ and $A$. We stress that this distribution is entirely determined by the $G(\cdot|X,A)$, so that this transformation only depends on the knowledge of both conditional survival functions $G$ and $S$, will be thus sometimes denoted $T^{*}_{\mathrm{DR}}(G,S)$ to underline this dependency. Importantly, this transformation not only satisfies the censoring-unbiasedness condition of Equation 14, but also enjoys a double robustness property: it remains valid as long as either $G$ or $S$ is correctly specified. The formal statement of this result is given in Proposition 9.6 in Appendix 9. The statement and proof of this results is found in Rubin and van der Laan, (2007) in the case where $C$ and $T$ are continuous.
@@ 3 Causal survival analysis in observational studies
Unlike RCT, observational data — such as from registries, electronic health records, or national healthcare systems — are collected without controlled randomized treatment allocation. In such cases, treated and control groups are likely unbalanced due to the non-randomized design, which results in the treatment effect being confounded by variables influencing both the survival outcome $T$ and the treatment allocation $A$. To enable identifiability of the causal effect, additional standard assumptions are needed.
@@
(Conditional Exchangeability / Unconfoundedness)
It holds
$A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X$
Under Assumption 3.1, the treatment assignment is randomly assigned conditionally on the covariates $X$. This assumption ensures that there are no unmeasured confounders as the latter would make it impossible to distinguish correlation from causality.
@@
(Positivity / Overlap for Treatment)
Letting $e(X):=\mathbb{P}(A=1\mid X)$ be the propensity score, there holds
$0<e(X)<1\quad\text{almost surely.}$
The Assumption 3.2 requires adequate overlap in covariate distributions between treatment groups, meaning every observation must have a non-zero probability of being treated. Because Assumption 2.1 (.Causal survival analysis in Randomized Controlled Trials) no longer holds anymore, neither does the previous identifiability Equation 4. In this new context, we can write
@@
$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X,A=1]-\mathbb{E}[T^{(0)}\wedge\tau|X,A=0]\right]$ (unconfoundness)
@@
In another direction, one could wish to identify the treatment effect through the survival curve as in Equation 1:
$S^{(a)}(t)=\mathbb{P}(T^{(a)}>t)=\mathbb{E}\left[\mathbb{P}(T>t|X,A=a)\right].$} (27)
Again, both identities still depend on the unknown quantity $T$ and suggest two different estimation strategies. These strategies differ according to the censoring assumptions and are detailed below.
@@ 3.1 Independent censoring
Figure 6 illustrates a causal graph in observational survival data with independent censoring (Assumptions 3.1 and 2.3).
@@
$S^{(1)}(t)=\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=1,T>t\}|X]}{\mathbb{E}[\mathbb{I}\{A=1\}|X]}\right]=\mathbb{E}\left[\frac{A\mathbb{I}\{T>t\}}{e(X)}\right],$
which suggests to adapt the classical KM estimator by reweighting it by the propensity score. The use of propensity score in causal inference has been initially introduced by Rosenbaum and Rubin, (1983) and further developed in Hirano et al., (2003). It was extended to survival analysis by Xie and Liu, (2005) through the adjusted Kaplan-Meier estimator as defined below.
@@
$D_{k}^{\mathrm{IPTW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{a}{\widehat{e}(X_{i})}+\frac{1-a}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
In Appendix 9, Proposition 9.7 establishes that $S^{*}_{\mathrm{IPTW}}(t\mid A=a)$ is a strongly consistent and asymptotically normal estimator of the survival function $S^{(a)}(t)$. Corollary 3.1 further shows that the same properties hold for the practical estimator provided the estimated propensity score converges uniformly. This follow directly from $S^{*}_{\mathrm{IPTW}}$ is a continuous function of $e(\cdot)$, and because $e$ and 1e1-e are lower-bounded as per Assumptions 3.2 (.Causal survival analysis in observational studies).
@@
which will be consistent under the same Assumptions as the previous Corollary. Note that we are not aware of any formal results concerning the bias and the asymptotic variance of the oracle estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$, and we refer to Xie and Liu, (2005) for heuristics concerning these questions.
@@ 3.2 Conditional independent censoring
Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.5 (.Conditionally independent censoring), the causal effect is affected both by confounding variables and by conditional censoring. The associated causal graph is depicted in Figure 7. In this setting, one can still use the $G$-formula exactly as in Section 2.2.1.
A natural alternative approach is to weight the IPCW and BJ transformations from Section 2.2.2 by the propensity score to disentangle both confounding effects and censoring at the same time.
@@
We mention that the G-formula approach developed in Section 2.2.1 does work in that context. In particular, Chen and Tsiatis, (2001) prove consistency and asymptotic normality results for Cox estimators in an observational study, and they give an explicit formulation of the asymptotic variance as a function of the parameters of the Cox model. In the non-parametric world, Foster et al., (2011) and Künzel et al., (2019) empirically study this estimator using survival forests, with the former employing it as a T-learner (referred to as Virtual Twins) and the latter as an S-learner.
@@ 3.2.1 IPTW-IPCW transformations
One can check that the IPCW transformation as introduced in Equation 18 is also a censoring-unbiased transformation in that context (formally stated in Proposition 9.8). Deriving an estimator of the difference in RMST is however slightly different in that context. In particular, Equation 15 rewrites
@@
Which in turn suggests to define
$\widehat{\theta}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{IPCW},i}.$} (30)
@@
Proposition 3.1.
@@
We can also use the same strategy as for the IPCW transformation and incorporate the new weights into a Kaplan-Meier estimator.
@@
$D_{k}^{\mathrm{IPTW-IPCW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW-IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(\widetilde{T}_{i}\wedge\tau|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
In Appendix 9, Proposition 9.9 shows that the oracle version of this estimator is consistent and asymptotically normal. The result extends to the practical estimator: as shown in the corollary below (Corollary 3.2), uniform consistency of the nuisance estimators ensures the consistency of the estimated survival curves $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}$.
@@
Under Assumptions 1.1 (.Definition of the estimand: the RMST), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the nuisance estimators $\widehat{G}(\cdot|A,X)$ and $\widehat{e}$ are weakly (resp. strongly) uniformly consistent, then $\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=a)$ is a weakly (resp. strongly) consistent estimator of $S^{(a)}(t)$.
We are not aware of general formulas for the asymptotic variances in this context. We mention nonetheless that Schaubel and Wei, (2011) have been able to derive asymptotic laws in this framework in the particular case of Cox-models.
@@ 3.2.2 IPTW-BJ transformations
Like IPCW transformation, BJ transformation is again a censoring-unbiased transformation in an observational context. This result is formalized in Proposition 9.10 in Appendix 9.
The corresponding estimator of the difference in RMST is
$\hat{\theta}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A}{e(X)}-\frac{1-A}{1-e(X)}\right)T^{*}_{\mathrm{BJ},i}.$} (32)
@@
Proposition 3.2.
@@ 3.2.3 Double augmented corrections
Building on the classical doubly-robust AIPTW estimator from causal inference (Robins et al.,, 1994), we could incorporate the doubly-robust transformations for censoring of Section 2.2.3 to obtain a doubly robust transformation for censoring and treatment assignment
@@
where we recall that $T^{*}_{\mathrm{DR}}$ is defined in Section 2.2.3. This transformation depends on four nuisance parameters: $G$ and $S$ through $T^{*}_{\mathrm{DR}}$, and now the propensity score $e$ and the conditional response $\mu$. This transformation does not really fall into the scope of censoring-unbiased transform, but it is easy to show that $\Delta^{*}_{\mathrm{QR}}$ is doubly robust for censoring and treatment assignment. That is, among the four nuisance functions involved, only one from each pair $(S,G)$ and $(e,\mu)$ needs to be correctly specified. This property (similar to Ozenne et al., (2020), Thm 1) is formally stated in Proposition 9.11 in Appendix 9.
@@
This estimator enjoys good asymptotic properties under parametric models, as detailed in Ozenne et al., (2020)
4 Variance estimation
Some survival estimators presented in this work have known asymptotic variance expressions, such as the Kaplan-Meier estimator (via Greenwood’s formula) or weighted Kaplan-Meier estimators. Under assumptions of asymptotic normality and independence between the two survival curves, the variance of the RMST difference can then be straightforwardly obtained through the delta method.
However, in practical situations, the covariance between estimated survival curves is typically nonzero, complicating variance estimation. While explicit variance formulas have been derived in certain settings—such as for the G-formula combined with semiparametric estimators (see Chen and Tsiatis, (2001))—these expressions often remain complex and difficult to implement. Therefore, we recommend using the nonparametric bootstrap (Efron and Tibshirani,, 1994) as a robust and practical solution for variance estimation and confidence interval construction in general settings. The effectiveness of bootstrap methods in the context of censored data has been investigated in several studies, including Efron, (1981) and Chen and Lo, (1996), which provide both theoretical justification and empirical validation for its use with Kaplan–Meier estimator.
5 Implementation
In this section, we first summarize the various estimators and their properties. We then provide custom implementations for all estimators, even those available in existing packages. These manual implementations serve two purposes: first, to make the methods accessible to the community when no existing implementation is available; and second, to facilitate a deeper understanding of the methods by detailing each step, even when a package solution exists. Finally, we present the packages available for directly computing $\theta_{\mathrm{RMST}}$.
5.1 Summary of the estimators
Table LABEL:tbl-nuisance provides an overview of the estimators introduced in this paper, along with the corresponding nuisance parameters needed for their estimation and an overview of their statistical properties in particular regarding their sensitivity to misspecification of the nuisance parameters.
Table 3: Estimators of the RMST difference and nuisance parameters needed to compute each estimator. Empty boxes indicate that the nuisance parameter is not needed in the estimator thus misspecification has no effect. Estimators in italic are those that are already implemented in available packages.

### v3

@@ 2.2.3 Augmented corrections
The main disadvantage of the two previous transformations is that they heavily rely on the specification of good estimator $\widehat{G}$ (for IPCW) or $\widehat{S}$ (for BJ). In order to circumvent this limitation, rubin2007 proposed the following transformations, whose expression is based on theory of semi-parametric estimation developed in vanderLaan2003,
@@
where $\mathop{}\!\mathrm{d}\mathbb{P}_{C}(t|X,A)$ is the distribution of $C$ conditional on the covariates $X$ and $A$. As this transformation depends on both conditional survival functions $G$ and $S$, we will denote it $T^{*}_{\mathrm{DR}}(G,S)$ to make this dependence explicit. Importantly, this transformation not only satisfies the censoring-unbiasedness condition of Equation 6, but also enjoys a double robustness property: it remains valid as long as either $G$ or $S$ is correctly specified. The formal statement of this result is given in Proposition 6.6 in Appendix A.1. Appendix A: Proofs. The statement and proof of this results is found in rubin2007 in the case where $C$ and $T$ are continuous.
@@ 3 Causal survival analysis in observational studies
Unlike RCT, observational data (such as from registries, electronic health records, or national healthcare systems) are collected without controlled randomized treatment allocation. In such cases, treated and control groups are likely unbalanced due to the non-randomized design, which results in the treatment effect being confounded by variables influencing both the survival outcome $T$ and the treatment allocation $A$. To enable identifiability of the causal effect, additional standard assumptions are needed.
@@
(Conditional Exchangeability / Unconfoundedness) It holds $A\perp\mkern-9.5mu\perp T^{(0)},T^{(1)}\mid X$
Under Assumption 3.1, the treatment assignment is randomly assigned conditionally on the covariates $X$. This assumption states that there are no unmeasured confounders as the latter would make it impossible to distinguish correlation from causality.
@@
(Positivity / Overlap for Treatment) Letting $e(X):=\mathbb{P}(A=1\mid X)$ be the propensity score, there holds $0<e(X)<1\quad\text{almost surely.}$
The Assumption 3.2 requires adequate overlap in covariate distributions between treatment groups, meaning every observation must have a non-zero probability of being treated and not being treated. Because Assumption 2.1 (.Causal survival analysis in Randomized Controlled Trials) no longer holds anymore, neither does the previous identifiability Equation 2. In this new context, we can write
@@
$=\mathbb{E}\left[\mathbb{E}[T^{(1)}\wedge\tau|X,A=1]-\mathbb{E}[T^{(0)}\wedge\tau|X,A=0]\right]$ (Unconfoundedness)
@@
A direct application of the above equation could be a direct plug-in G-formula estimator developed in Section 2.2.1, equation 4. In particular, chen2001 prove consistency and asymptotic normality results for Cox estimators in an observational study, and they give an explicit formulation of the asymptotic variance as a function of the parameters of the Cox model. In the non-parametric literature, foster2011 and kunzel2019 empirically study this estimator using survival forests, with the former employing it as a T-learner (referred to as Virtual Twins) and the latter as an S-learner.
In another direction, the next sections describe alternative estimation strategies that also rely on equation 16 but adapt the estimation of the conditional expectations to the specific censoring assumptions—either independent or conditionally independent censoring.
@@ 3.1 Independent censoring
Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.3 (.Independent censoring: the Kaplan-Meier estimator), the data exhibit confounding effect induced by the confounders. The corresponding causal graph is shown in Figure 5.
@@
$\theta_{\mathrm{RMST}}$ $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau|X,A=1]-\mathbb{E}[T\wedge\tau|X,A=0]\right]$ (17)
$=\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=1\}T\wedge\tau|X]}{\mathbb{E}[\mathbb{I}\{A=1\}|X]}\right]-\mathbb{E}\left[\frac{\mathbb{E}[\mathbb{I}\{A=0\}T\wedge\tau|X]}{\mathbb{E}[\mathbb{I}\{A=0\}|X]}\right]$ (Positivity for treatment)
$=\mathbb{E}\left[\frac{AT\wedge\tau}{e(X)}\right]-\mathbb{E}\left[\frac{(1-A)T\wedge\tau}{1-e(X)}\right],$
which suggests to adapt the classical KM estimator by reweighting it by the propensity score (see equation 1). The use of propensity score in causal inference has been initially introduced by rosenbaum1983 and further developed in hirano2003. It was extended to survival analysis by xie2005 through the adjusted Kaplan-Meier estimator as defined below.
@@
$D_{k}^{\mathrm{IPTW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}=t_{k},\Delta_{i}=1,A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
In Appendix A.1. Appendix A: Proofs, Proposition 6.7 establishes that $S^{*}_{\mathrm{IPTW}}(t\mid A=a)$ is a strongly consistent and asymptotically normal estimator of the survival function $S^{(a)}(t)$. Corollary 3.1 further shows that the same properties hold for the practical estimator provided the estimated propensity score converges uniformly. This follow directly from $S^{*}_{\mathrm{IPTW}}$ is a continuous function of $e(\cdot)$, and because $e$ and 1e1-e are lower-bounded as per Assumptions 3.2 (.Causal survival analysis in observational studies).
@@
which will be consistent under the same Assumptions as the previous Corollary. Note that we are not aware of any formal results concerning the bias and the asymptotic variance of the oracle estimator $S^{*}_{\mathrm{IPTW}}(t|A=a)$, and we refer to Xie and Liu (2005) for heuristics concerning these questions.
@@ 3.2 Conditional independent censoring
Under Assumptions 3.1 (.Causal survival analysis in observational studies) and 2.5 (.Conditionally independent censoring), the data exhibit both confounding and conditionally independent censoring. The corresponding causal graph is shown in Figure 6. To address this, weight the IPCW and Buckley–James (BJ) transformations from Section 2.2.2 by the propensity score.
@@ 3.2.1 IPTW-IPCW transformations
One can check that the IPCW transformation as introduced in Equation 10 is also a censoring-unbiased transformation in observational study (see Appendix in Proposition 6.8). Deriving an estimator of the difference of RMST is however slightly different in that context. In particular, Equation 7 rewrites
@@
which suggests defining an estimator of the RMST difference that can, similarly to IPCW estimators, be stabilized by normalizing with the sum of the weights, and is identified by combining Equation 17 with the IPCW representation in Equation 10.
$\widehat{\theta}_{\mathrm{IPTW-IPCW}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)T^{*}_{\mathrm{IPCW},i}.$} (20)
@@
Proposition 3.1 (Consistency of $\widehat{\theta}_{\mathrm{IPTW-IPCW}}$ with estimated $\hat{e}$ and $\hat{G}$).
@@
We can also use the same strategy as for the IPCW transformation and incorporate the time-dependent weights into a Kaplan-Meier estimator.
@@
$D_{k}^{\mathrm{IPTW-IPCW}}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{\Delta_{i}^{\tau}}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}=t_{k},A_{i}=a)$
$\quad\text{and}\quad N^{\mathrm{IPTW-IPCW}}_{k}(a)$ $:=\sum_{i=1}^{n}\left(\frac{A_{i}}{\widehat{e}(X_{i})}+\frac{1-A_{i}}{1-\widehat{e}(X_{i})}\right)\frac{1}{\widehat{G}(t_{k}|X_{i},A=a)}\mathbb{I}(\widetilde{T}_{i}\geqslant t_{k},A_{i}=a),$
@@
In Appendix A.1. Appendix A: Proofs, Proposition 6.9 shows that the oracle version of this estimator is consistent and asymptotically normal. The result extends to the empirical estimator: as shown in Corollary 3.2, uniform consistency of the nuisance estimators ensures the consistency of the estimated survival curves $\widehat{\theta}_{\mathrm{IPTW-IPCW-KM}}$.
@@
Under Assumptions 1.1 (.Notations and definition of the Restricted Mean Survival Time), 3.1 (.Causal survival analysis in observational studies), 3.2 (.Causal survival analysis in observational studies), 2.5 (.Conditionally independent censoring) and 2.6 (.Conditionally independent censoring), and for all $t\in[0,\tau]$, if the nuisance estimators $\widehat{G}(\cdot|A,X)$ and $\widehat{e}$ are weakly (resp. strongly) uniformly consistent, then $\widehat{S}_{\mathrm{IPTW-IPCW}}(t|A=a)$ is a weakly (resp. strongly) consistent estimator of $S^{(a)}(t)$.
We are not aware of general formulas for the asymptotic variances in this context. We mention nonetheless that schaubel2011 have been able to derive asymptotic laws in this framework in the particular case of Cox-models.
@@ 3.2.2 IPTW-BJ transformations
As with IPCW, the BJ transformation in Equation 13 is censoring-unbiased in observational contexts (Proposition 6.10, Appendix A.1. Appendix A: Proofs). The corresponding estimator of the difference in RMST is
$\hat{\theta}_{\mathrm{IPTW-BJ}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(X_{i})}-\frac{1-A_{i}}{1-e(X_{i})}\right)T^{*}_{\mathrm{BJ},i}.$} (22)
@@
Proposition 3.2 (Consistency of $\widehat{\theta}_{\mathrm{IPTW-BJ}}$ with estimated $\hat{e}$ and $\hat{S}$).
@@ 3.2.3 Double augmented corrections
Building on Equations 16 and 17, it follows that
$\mathbb{E}\!\left[\frac{\mathbb{I}\{A=a\}}{\mathbb{E}[\mathbb{I}\{A=a\}\mid X]}\,\Big((T\wedge\tau)-\mathbb{E}[T\wedge\tau\mid X,A=a]\Big)\,\middle|\,X\right]=0.$
Thus,
$\theta_{\mathrm{RMST}}$ $=\mathbb{E}\left[\mathbb{E}[T\wedge\tau\mid X,A=1]-\mathbb{E}[T\wedge\tau\mid X,A=0]\right]$ (23)
$=\mathbb{E}\left[\mathbb{E}[T\wedge\tau\mid X,A=1]-\mathbb{E}[T\wedge\tau\mid X,A=0]\right$.
$\qquad\left.+\underbrace{\frac{\mathbb{I}\{A=1\}}{e(X)}\Big(T\wedge\tau-\mathbb{E}[T\wedge\tau\mid X,A=1]\Big)-\frac{\mathbb{I}\{A=0\}}{1-e(X)}\Big(T\wedge\tau-\mathbb{E}[T\wedge\tau\mid X,A=0]\Big)}_{\text{\small add mean-zero terms}}\right].$
This representation coincides with the classical doubly robust augmented inverse probability weighting (AIPTW) functional from causal inference (robins1994). Incorporating the doubly robust censoring transformations from Section 2.2.3 yields a transformation that is doubly robust with respect to both censoring and treatment assignment
@@
where we recall that $T^{*}_{\mathrm{DR}}$ is defined in Section 2.2.3. This transformation depends on four nuisance parameters: $G$ and $S$ through $T^{*}_{\mathrm{DR}}$, and now the propensity score $e$ and the conditional response $\mu$. This transformation does not really fall into the scope of censoring-unbiased transform, but it is easy to show that $\Delta^{*}_{\mathrm{QR}}$ is doubly robust for censoring and treatment assignment. That is, among the four nuisance functions involved, only one from each pair $(S,G)$ and $(e,\mu)$ needs to be correctly specified. This property (similar to Ozenne20_AIPTW_AIPCW, Thm 1) is formally stated in Proposition 6.11 in Appendix A.1. Appendix A: Proofs.
@@
Asymptotic properties of this estimator, such as asymptotic normality and semiparametric efficiency, are detailed in Ozenne20_AIPTW_AIPCW.
4 Summary of the estimators and variance estimation
For completeness, Table LABEL:tbl-nuisance offers a compact synopsis of the estimators, context of application, associated nuisance models, and large-sample behavior, with a focus on misspecification sensitivity.
Table 3: Estimators of the RMST difference and nuisance parameters needed to compute each estimator. Empty boxes indicate that the nuisance parameter is not needed in the estimator thus misspecification has no effect.